OpenStudy (anonymous):

A 0.400kg ice puck, moving east with a speed of 5.24m/s, has a head-on collision with a 0.900kg puck initially at rest. Assume the collision is perfectly elastic. What is the speed of the 0.400 puck after the collision?

OpenStudy (anonymous):

So far, I've plugged it into the equation V1i(m1-m2/m1+m2)=2V1f and I came out with 16.5

OpenStudy (anonymous):

is that the equation that they gave you or did you make that yourself?

OpenStudy (anonymous):

I just made it up myself. I'm at a lose with this problem.

OpenStudy (caominhim):

\[V_{1} = \frac{M_{1}-M_{2}}{M_{1}+M_{2}}V_{o1} - \frac{2M_{2}}{M_{1}+M_{2}}V_{o2}\] in your case Vo2 is 0